# Embodied AI in Urban Spaces

### v1

@@ 3.2 Scalable Object Retrieval
Hierarchical layout generation decides the scene’s layout and where to place the objects. However, to make the trained agents generalizable when navigating through scenes composed of various objects in the real world, what objects to place is another crucial question. In this section, we propose the Scalable Assets Retrieval pipeline, in which we first get real-world object distributions from web data, and then retrieve objects from 3D asset repositories through an open-vocabulary search schema based on VLMs. This pipeline is flexible and extensible: the retrieved objects can be scaled to arbitrary sizes as we continue to exploit more web data for scene descriptions and include more 3D assets as the candidate objects.
@@
Real-world object distribution extraction. Urban spaces have unique structures and object distributions, such as the infrastructure built by the urban planning administration and clutters placed by people. Thus, we design a real-world distribution extraction method to get a description pool depicting the frequent objects in urban spaces. As illustrated in Figure 5 (a), we first leverage off-the-shelf academic datasets for scene understanding, CityScape [10] and Mapillary Vistas [52], to get a list of 90 objects that are with high frequency to be put in the urban space. However, the number of objects is limited because of the closed-set definitions in the image datasets. We introduce two open-set sources to get broader object distribution from the real world. 1) Google Street data. We first collect 25,000 urban space images from 50 countries across six continents. Then, we harness GPT-4o [53] and open-set segmentation model Grounded-SAM [66] to get 1,075 descriptions of objects in the urban public space. 2) Urban planning description data. We further get a list of 50 essential objects in public urban spaces through a thorough survey of 10 urban design handbooks. Finally, by combining these three data sources, we can get an object description pool with 1,215 items of descriptions that form the real-world object category distribution.
Figure 5: Scalable assets retrieval. (a) Real-world distribution extraction. We get object distribution for urban spaces from three sources: academic datasets, Google Street data, and text description data. (b) Open-vocabulary search. We use the VLM to get image and text embedding, respectively. Then, based on the relevant scores, we can get the objects with high rankings.
@@
Open-vocabulary search. The recent development of large-scale 3D object repositories [16, 14, 80] enables efficiently constructing a dataset for a specific scene. However, these large repositories have three intrinsic issues to harness these repositories: 1) most of the data is unrelated to the urban scene, 2) the data quality in large repositories is uneven, and 3) the data has no reliable attribute annotations. To this end, we introduce an open-vocabulary search method to tackle these issues. As shown in Figure 5 (b), the whole pipeline is based on an image-text retrieval architecture. We first sample objects from Objaverse [16] and Objaverse-XL [14] repositories to get projected multi-view images. Here, a naive uniform view sampling will bring low-quality harmful images. Following [49, 48], we select and prioritize informative viewpoints, which significantly enhance retrieval effectiveness. Then, we leverage the encoder of a Vision Language Model BLIP [43] to extract features from projected images and sampled descriptions from the object description pool, respectively, to calculate relevant scores. Then, we can get target objects with relevant scores up to a threshold. This method lets us get an urban-specific dataset with 10,000 high-quality objects in real-world category distributions. In addition, we provide an interface for customizing training objects in the scene by providing images or text descriptions, taking advantage of recent advances in 3D object reconstruction [46, 34] and generation [60, 9].
@@ 3.3 Cohabitant Populating
Following BEDLAM [6] and AGORA [57], we represent humans as parametric human model SMPL-X [58], in which the 3D human body is controlled by a set of parameters for pose $\theta$, shape $\beta$, and facial expression $\phi$, respectively. Then, built upon SynBody [81]’s asset repository, 1,100 3D rigged human models are procedually generated by sampling from 68 garments, 32 hairs, 13 beards, 46 accessories, and 1,038 cloth and skin textures. To form safety-critical scenarios, we also include vulnerable road users like bikers, skateboarders, and scooter riders. For the other agents, we incorporate the 3D assets of COCO Robotics and Starship’s delivery robots, Drive Medical’s electric wheelchair, Boston Dynamic’s robot dog, and Agility Robotics’ humanoid robot.
We provide two kinds of human movements in the simulator – daily movements and unique movements. Daily movements provide the basic human dynamics in daily life, i.e., idle, walking, and running. Unique movements are the complicated dynamics that appear randomly in public spaces, such as dancing and exercising. We harness the BEDLAM dataset [6] to obtain 2,311 unique movements. For humans and other agents with daily movements, we simulate their trajectories using the ORCA [77] social forces model and Push and Rotate (P&R) algorithm [11]. ORCA [77] uses a joint optimization and a centralized controller that guarantees that agents will not collide with each other or any other objects identified as obstacles. Push and Rotate (P&R) [11] is a multi-agent path-finding algorithm that can resolve any potential deadlock by local coordination. In the future, an interesting direction is to endow personal traits like job, personality, and purpose to humans and harness the advantages of LLMs [1] and LVMs [45] to enable social [62] and interactive behaviors [55] of humans in urban scenes.
4 MetaUrban-12K Dataset
@@
Figure 6: MetaUrban-12K statistics.
@@
5 Experiments
Experimental settings.
@@ 5 Experiments
Tasks. We design two common tasks in urban scenes: Point Navigation (PointNav) and Social Navigation (SocialNav). In PointNav, the agent’s goal is to navigate to the target coordinates in static environments without access to a pre-built environment map. In SocialNav, the agent is required to reach a point goal in dynamic environments that contain moving environmental agents. The agent shall avoid collisions or proximity to environmental agents beyond thresholds to avoid penalization (distance <0.2 meters). The agent’s action space in the experiments consists of acceleration, brake, and steering. The observations contain a vector denoting the LiDAR signal, a vector summarizing the agent’s state, and the navigation information that guides the agent toward the destination. Methods. We evaluate 7 typical baseline models to build comprehensive benchmarks on MetaUrban, across Reinforcement Learning (PPO [68]), Safe Reinforcement Learning (PPO-Lag [64], and PPO-ET [72]), Offline Reinforcement Learning (IQL [36] and TD3+BC [23]), and Imitation Learning (BC [3] and GAIL [30]). Evaluation metrics. The agent is evaluated using the Success Rate (SR) and Success weighted by Path Length (SPL) [2, 4] metrics, which measure the success and efficiency of the path taken by the agent. For SocialNav, except Success Rate (SR), the Social Navigation Score (SNS) [12], is also used to evaluate the social complicity of the agent. For both tasks, we further report the Cumulative Cost (CC) [44] to evaluate the safety properties of the agent. It records the crash frequency to obstacles or environmental agents. Please refer to the Appendix for details of models, rewards, and hyperparameters.
Table 1: Benchmarks. The benchmark of PointNav and SocialNav tasks on the MetaUrban-12K dataset. Seven representative methods of RL, safe RL, offline RL, and IL are evaluated for each benchmark.     indicate the best performance among online methods (RL and Safe RL) and offline methods (offline RL and IL), respectively.
Category Method PointNav SocialNav
Test Zero-shot Test Zero-shot
@@
RL PPO [68] 0.66 0.64 0.51 0.49 0.45 0.78 0.34 0.64 0.66 0.24 0.57 0.51
Safe RL PPO-Lag [64] 0.60 0.58 0.41 0.60 0.57 0.53 0.17 0.51 0.33 0.08 0.47 0.50
PPO-ET [72] 0.57 0.53 0.47 0.53 0.49 0.65 0.05 0.52 0.26 0.02 0.50 0.62
Offline RL IQL [36] 0.36 0.33 0.49 0.30 0.27 0.63 0.36 0.67 0.39 0.27 0.62 3.05
TD3+BC [23] 0.29 0.28 0.77 0.20 0.20 1.16 0.26 0.61 0.62 0.32 0.64 1.53
IL BC [3] 0.36 0.28 0.83 0.32 0.26 1.15 0.28 0.56 1.23 0.18 0.54 0.58
GAIL [30] 0.47 0.36 1.05 0.40 0.32 1.46 0.34 0.63 0.71 0.28 0.61 0.67
5.1 Benchmarks
We build two benchmarks on the MetaUrban-12K dataset for PointNav and SocialNav tasks. We train 7 typical baselines on the MetaUrban-train dataset and then evaluate them on the MetaUrban-test set. We use the demonstration data provided in MetaUrban-12K for offline RL and IL training. We further make zero-shot evaluations on the MetaUrban-unseen set to demonstrate the generalizability of models trained on the MetaUrban-12K dataset while directly tested on unseen environments.
Table 1 shows the results in the PointNav and SocialNav benchmarks. From the results, we can draw 4 key observations. 1) The tasks are far from being solved. The highest success rates are only $66\%$ and $36\%$ for PointNav and SocialNav tasks achieved by the baselines, indicating the difficulty of these tasks in the urban environments composed by MetaUrban. Note that these benchmarks are built on a medium level of object and dynamic density; increasing the density will further degrade the performances shown in ablation studies. 2) Models trained on MetaUrban-12K have strong generalizability in unseen environments. With zero-shot testing, models can still achieve $41\%$ and $26\%$ success rates on average for PointNav and SocialNav tasks. These results are strong since the models generalize to not only unseen objects and layouts but also unseen dynamics of environmental agents. It demonstrates that the compositional nature of MetaUrban, supporting the coverage of a large spectrum of complex urban scenes, can successfully empower generalization ability to the trained models. 3) SocialNav is much harder than PointNav due to the dynamics of the mobile environmental agents. On average, the success rate decreases by $15\%$ from PointNav to SocialNav, indicating that dynamic agents, such as pedestrians, vulnerable road users, and other agents common in urban scenes, present significant challenges to the trained agent. 4) Safe RL remarkably improves the safety property at the expense of effectiveness. Among all tasks and settings, the safe RL models achieve the best performance in the Cumulative Cost, indicating that these models are successful at avoiding collision with pedestrians and objects. However, the success rate and SPL or SNS decrease accordingly, indicating future efforts to balance the safety and effectiveness of agents in complex urban scenes.
@@ 5.2 Ablation Study
Evaluation of generalizability. To evaluate the generalizable ability of agents trained on data generated by MetaUrban, we compare the success rate of four settings in Figure 7 (a). Setting-1 and Setting-2 are the results of training on MetaUrban-train while testing on MetaUrban-test and MetaUrban-unseen, respectively. We can observe a performance drop on MetaUrban-unseen. However, the zero-shot results still achieve $49\%$ success rate facing various out-of-distribution scenes, demonstrating the strong generalizability of models trained on large-scale data created by MetaUrban. Setting-3 and Setting-4 are the results of direct training on MetaUrban-finetune, and fine-tuning on MetaUrban-finetune from the pre-trained model on MetaUrban-train. Compared between Setting-2 and Setting-3, we can observe an obvious performance drop, which is caused by an underfitting of the insufficient and complex fine-tuning data. Setting-4 outperforms Setting-3 by a large margin, demonstrating that the model trained on the MetaUrban-12K dataset can provide informative priors as good initializations for quick tuning.
@@ 6 Conclusion
We propose a new compositional simulator, MetaUrban, to facilitate embodied AI and robotics research in urban scenes. MetaUrban can generate infinite urban environments with complex scene structures and diverse movements of pedestrians and other mobile agents. These environments used as training data can significantly improve the generalizability and safety of the embodied AI underlying different mobile machines from food delivery bots to humanoids. We commit ourselves to developing the open-source simulator and fostering the community effort to turn it into a sustainable community infrastructure.

### v2

@@ 3.2 Scalable Obstacle Retrieval
Hierarchical layout generation decides the scene’s layout and where to place the objects. However, to make the trained agents generalizable when navigating through scenes composed of various objects in the real world, what objects to place is another crucial question. In this section, we propose the Scalable Obstacle Retrieval pipeline, in which we first get real-world object distributions from web data, and then retrieve objects from 3D asset repositories through an open-vocabulary search schema based on VLMs. This pipeline is flexible and extensible: the retrieved objects can be scaled to arbitrary sizes as we continue to exploit more web data for scene descriptions and include more 3D assets as the candidate objects.
@@
Real-world object distribution extraction. Urban spaces have unique structures and object distributions, such as the infrastructure built by the urban planning administration and clutters placed by people. Thus, we design a real-world distribution extraction method to get a description pool depicting the frequent objects in urban spaces. As illustrated in Figure 5 (a), we first leverage off-the-shelf academic datasets for scene understanding, CityScape (Cordts et al., 2016) and Mapillary Vistas (Neuhold et al., 2017), to get a list of 90 objects that are with high frequency to be put in the urban space. However, the number of objects is limited because of the closed-set definitions in the image datasets. We introduce two open-set sources to get broader object distribution from the real world. 1) Google Street data. We first collect 25,000 urban space images from 50 countries across six continents. Then, we harness GPT-4o (OpenAI, 2024) and open-set segmentation model Grounded-SAM (Ren et al., 2024) to get 1,075 descriptions of objects in the public urban space. 2) Urban planning description data. We further get a list of 50 essential objects in public urban spaces through a thorough survey of 10 urban design handbooks. Finally, by combining these three data sources, we can get an object description pool with 1,215 items of descriptions that form the real-world object category distribution.
@@
Open-vocabulary search. The recent development of large-scale 3D object repositories (Deitke et al., 2023b; 2024; Wu et al., 2023) enables efficiently constructing a dataset for a specific scene. However, these large repositories have three intrinsic issues to harness these repositories: 1) most of the data is unrelated to the urban scene, 2) the data quality in large repositories is uneven, and 3) the data has no reliable attribute annotations. To this end, we introduce an open-vocabulary search method to tackle these issues. As shown in Figure 5 (b), the whole pipeline is based on an image-text retrieval architecture. We first sample objects from Objaverse (Deitke et al., 2023b) and Objaverse-XL (Deitke et al., 2024) repositories to get projected multi-view images. Here, a naive uniform view sampling will bring low-quality harmful images. Following (Luo et al., 2023; 2024), we select and prioritize informative viewpoints, which significantly enhance retrieval effectiveness. Then, we leverage the encoder of a Vision Language Model BLIP (Li et al., 2022a) to extract features from projected images and sampled descriptions from the object description pool, respectively, to calculate relevant scores. Then, we can get target objects with relevant scores up to a threshold. This method lets us get an urban-specific dataset with 10,000 high-quality obstacles in real-world category distributions.
Figure 5: Scalable obstacle retrieval. (a) Real-world distribution extraction. We get object distribution for urban spaces from three sources: academic datasets, Google Street data, and text description data. (b) Open-vocabulary search. We use the VLM to get image and text embedding, respectively. Then, based on the relevant scores, we can get the objects with high rankings.
@@ 3.3 Cohabitant Populating
Following BEDLAM Black et al. (2023) and AGORA Patel et al. (2021), we represent humans as parametric human model SMPL-X Pavlakos et al. (2019), in which the 3D human body is controlled by a set of parameters for pose $\theta$, shape $\beta$, and facial expression $\phi$, respectively. Then, built upon SynBody Yang et al. (2023)’s asset repository, 1,100 3D rigged human models are procedually generated by sampling from 68 garments, 32 hairs, 13 beards, 46 accessories, and 1,038 cloth and skin textures. To form safety-critical scenarios, we also include vulnerable road users like bikers, skateboarders, and scooter riders. For the other agents, we incorporate the 3D assets of COCO Robotics’, Starship’s and Yandex’s delivery bots, Drive Medical’s electric wheelchair, Pride Go-Go’s mobility scooter, Boston Dynamic’s robot dog, and Agility Robotics’ humanoid robot.
We provide two kinds of human movements in the simulator – daily movements and unique movements. Daily movements provide the basic human dynamics in daily life, i.e., idle, walking, and running. Unique movements are the complicated dynamics that appear randomly in public spaces, such as dancing and exercising. We harness the BEDLAM dataset Black et al. (2023) to obtain 2,311 unique movements. For humans and other agents with daily movements, we simulate their trajectories using the ORCA Van Den Berg et al. (2011) social forces model and Push and Rotate (P&R) algorithm De Wilde et al. (2014). ORCA Van Den Berg et al. (2011) uses a joint optimization and a centralized controller that guarantees that agents will not collide with each other or any other objects identified as obstacles. Push and Rotate (P&R) De Wilde et al. (2014) is a multi-agent path-finding algorithm that can resolve any potential deadlock by local coordination. In the future, an interesting direction is to endow personal traits like job, personality, and purpose to humans and harness the advantages of LLMs Achiam et al. (2023) and LVMs Liu et al. (2023a) to enable social Puig et al. (2023a) and interactive behaviors Park et al. (2023) of humans in urban scenes.
4 Experiments
In this section, we first introduce the data and evaluation metrics used in all the experiments. In Section 4.1, we build comprehensive benchmarks for two core tasks in micromobility – point navigation and social navigation, across seven typical baseline models. In Section 4.2, we make evaluations across different mobile machines to delve into the influence of mechanical structures, such as engine force, wheel friction, and wheelbase, on their capability when learning and executing policies. In Section 4.3, we evaluate the generalizability and scaling ability of the MetaUrban platform, and reveal the effects of the density of static objects and dynamic agents.
Data.
@@ 4 Experiments
Evaluation metrics.
The agent is evaluated using the Success Rate (SR) and Success weighted by Path Length (SPL) (Anderson et al., 2018; Batra et al., 2020) metrics, which measure the success and efficiency of the path taken by the agent. For SocialNav, except Success Rate (SR), the Social Navigation Score (SNS) (Deitke et al., 2022a), is also used to evaluate the social complicity of the agent. For both tasks, we further report the Cumulative Cost (CC) (Li et al., 2022b) to evaluate the safety properties of the agent. It records the crash frequency to obstacles or environmental agents.
Table 1: Benchmarks. The benchmark of PointNav and SocialNav tasks on the MetaUrban-12K dataset. Seven representative methods of RL, safe RL, offline RL, and IL are evaluated for each benchmark.   indicate the best performance among online methods (RL and Safe RL).   indicates the best performance among offline methods (offline RL and IL).
Category Method PointNav SocialNav
Test Unseen Test Unseen
@@
RL PPO (Schulman et al., 2017) 66% 0.64 0.51 49% 0.45 0.78 34% 0.64 0.66 24% 0.57 0.51
Safe RL PPO-Lag (Ray et al., 2019) 60% 0.58 0.41 60% 0.57 0.53 17% 0.51 0.33 8% 0.47 0.50
PPO-ET (Sun et al., 2021) 57% 0.53 0.47 53% 0.49 0.65 5% 0.52 0.26 2% 0.50 0.62
Offline RL IQL (Kostrikov et al., 2021) 36% 0.33 0.49 30% 0.27 0.63 36% 0.67 0.39 27% 0.62 3.05
TD3+BC (Fujimoto & Gu, 2021) 29% 0.28 0.77 20% 0.20 1.16 26% 0.61 0.62 32% 0.64 1.53
IL BC (Bain & Sammut, 1995) 36% 0.28 0.83 32% 0.26 1.15 28% 0.56 1.23 18% 0.54 0.58
GAIL (Ho & Ermon, 2016) 47% 0.36 1.05 40% 0.32 1.46 34% 0.63 0.71 28% 0.61 0.67
4.1 Benchmarks
In micromobility, the core and foundational function for a mobile machine is to navigate from point A to point B within a bustling urban environment (Mitchell et al., 2010; Gössling, 2020). In this function, there are two main commands: 1) not collide with static objects (infrastructures and clutters), and 2) not bump into moving objects (pedestrians and other agents). To this end, we design two foundational tasks – point navigation and social navigation, to evaluate the generalizability and safety of recent state-of-the-art embodied AI models on a typical wheeled mobile machine4.1, i.e., COCO Robotics Delivery Bot4.1.
33footnotetext: Note that MetaUrban provides an easy-to-use interface for users to benchmark the navigation ability of any mobile machine they want to evaluate, such as electric wheelchairs, robot dogs, and humanoids. In a navigation-locomotion framework (Lee et al., 2024), an off-the-shelf locomotion module can be fixed for one machine; users can change the navigation module independently to benchmark.44footnotetext: https://www.cocodelivery.com/
Tasks.
@@ 4.1 Benchmarks
We design two common tasks in urban scenes: Point Navigation (PointNav) and Social Navigation (SocialNav). In PointNav, the agent’s goal is to navigate to the target coordinates in static environments without access to a pre-built environment map. In SocialNav, the agent is required to reach a point goal in dynamic environments that contain moving environmental agents. The agent shall avoid collisions or proximity to environmental agents beyond thresholds to avoid penalization (distance <0.2 meters). The agent’s action space in the experiments consists of acceleration, brake, and steering. The observations contain a vector denoting the LiDAR signal, a vector summarizing the agent’s state, and the navigation information that guides the agent toward the destination.
Models.
We build two benchmarks on the MetaUrban-12K dataset for PointNav and SocialNav tasks. We evaluate 7 typical baseline models, across Reinforcement Learning (PPO (Schulman et al., 2017)), Safe Reinforcement Learning (PPO-Lag (Ray et al., 2019), and PPO-ET (Sun et al., 2021)), Offline Reinforcement Learning (IQL (Kostrikov et al., 2021) and TD3+BC (Fujimoto & Gu, 2021)), and Imitation Learning (BC (Bain & Sammut, 1995) and GAIL (Ho & Ermon, 2016)). We train all these baseline models on the MetaUrban-train dataset and then evaluate them on the MetaUrban-test set. We use the demonstration data provided in MetaUrban-12K for offline RL and IL training. We further make unseen evaluations on the MetaUrban-unseen set to demonstrate the generalizability of models trained on the MetaUrban-12K dataset while directly tested on unseen environments. Please refer to Appendix D for details of models, rewards, and hyperparameters.
Results.
Table 1 shows the results in the PointNav and SocialNav benchmarks. We can draw two key observations from the results. 1) The tasks are far from being solved. The highest success rates are only $66\%$ and $36\%$ for PointNav and SocialNav tasks achieved by the baselines, indicating the difficulty of these tasks in the urban environments composed by MetaUrban. Note that these benchmarks are built on a medium level of object and dynamic density; increasing the density will further degrade the performances shown in ablation studies. 2) Models trained on MetaUrban-12K have strong generalizability in unseen environments. With unseen evaluation, models can still achieve $41\%$ and $26\%$ success rates on average for PointNav and SocialNav tasks. These results are strong since the models generalize to not only unseen objects and layouts but also unseen dynamics of environmental agents. It demonstrates that the compositional nature of MetaUrban, supporting the coverage of a large spectrum of complex urban scenes, can successfully empower generalization ability to the trained models. In addition, other interesting observations include: SocialNav is much harder than PointNav due to the dynamics of the mobile environmental agents, and Safe RL remarkably improves the safety properties at the expense of effectiveness.
4.2 Evaluation across Mobile Machines
Different mobile machines have heterogeneous mechanical structure design specifications, which have a huge influence on their navigation behaviors and capabilities. MetaUrban can evaluate different mechanical structures and designs of mobile machines before deployment and volume production. In this section, we evaluate three typical wheeled mobile machines (a delivery bot, an electric wheelchair, and a mobility scooter) with large design variations, and investigate the relationship between mechanical structures and performance in policy learning. Further, we also investigate the influence of mechanical structures in policy execution on different terrains, as detailed in Appendix D.3.
Table 2: Evaluation of policy learning across mobile machines.   and   indicate the best and worst performance among machines in the straight street block (“S”).   and   indicate the best and worst performance among machines in the intersection street block (“X”).
COCO (base) COCO (mod-1) COCO (mod-2) Wheelchair Mobility Scooter
Max speed $v_{max}$ ($km/h$) 30 10 30 5 45
Max steering $s_{max}$ ($degree$) 40 40 10 5 45
Wheel friction $\mu$ 0.7 0.7 0.7 0.7 0.7
Engine force $F$ ($N$) $350\sim 550$ $350\sim 550$ $350\sim 550$ $100\sim 200$ $450\sim 650$
Brake force $B$ ($N$) $35\sim 80$ $35\sim 80$ $35\sim 80$ $35\sim 80$ $35\sim 80$
@@ 4.2 Evaluation across Mobile Machines
[…]
In this experiment, we evaluate the influence of mechanical structures on policy learning in a static obstacle avoidance task. We follow the setting of PointNav experiments in Section 1. We perform PPO model training on three mobile machines, having significant disparities in max speed, max steering, wheel friction, engine force, and brake force. In addition, apart from comparing different machines, we compare different variations of one machine, i.e., COCO (base), COCO (mod-1), and COCO (mod-2), to find a better specification that can satisfy the demands of different use cases. We evaluate all five models in straight street blocks (mainly static obstacles) and intersection street blocks (dense interactions with pedestrians) to test their behaviors in different scenarios, which are noted as “S” and “X” in Table 2. We use “Success Rate (SR)” and “Success weighted by Path Length (SPL)” to measure the mobility of agents, while use “Cost” to measure the safety of agents.
[…]
@@ 4.3 Ablation Study
Evaluation of generalizability. To evaluate the generalizable ability of agents trained on data generated by MetaUrban, we compare the success rate of four settings in Figure 6 (a). Setting-1 and Setting-2 are the results of training on MetaUrban-train while testing on MetaUrban-test and MetaUrban-unseen, respectively. We can observe a performance drop on MetaUrban-unseen. However, the unseen evaluation results still achieve $49\%$ success rate facing various out-of-distribution scenes, demonstrating the strong generalizability of models trained on large-scale data created by MetaUrban. Setting-3 and Setting-4 are the results of direct training on MetaUrban-finetune, and fine-tuning on MetaUrban-finetune from the pre-trained model on MetaUrban-train. Compared between Setting-2 and Setting-3, we can observe an obvious performance drop, which is caused by an underfitting of the insufficient and complex fine-tuning data. Setting-4 outperforms Setting-3 by a large margin, demonstrating that the model trained on the MetaUrban-12K dataset can provide informative priors as good initializations for quick tuning.
@@ 5 Conclusion
We propose a new compositional simulator, MetaUrban, to facilitate embodied AI and micromobility research in urban scenes. MetaUrban can generate infinite interactive urban environments with complex scenes, diverse obstacles, and diverse movements of pedestrians and other mobile agents. These environments used as training data can significantly improve the generalizability and safety of the embodied AI underlying mobile machines. We commit ourselves to developing the open-source simulator and fostering community efforts to turn it into a sustainable infrastructure.
@@
In the appendix, we present more details of MetaUrban. In Section A, we illustrate samples of static and dynamic scenes, as well as static and dynamic assets in MetaUrban. In Section B, we elaborate on the technical details of the MetaUrban simulator. In Section C, we will give the construction details and statistics of the MetaUrban-12K dataset. In Section D, we will present details of implementations in our benchmarks and further exploration in cross-machine evaluations. In Section E, we delve into and validate unique challenges in urban micromobility. In Section F, we provide the datasheet of MetaUrban and MetaUrban-12K dataset. In Section G, and  H, we evaluate the performance and robustness of models trained on MetaUrban. In Section I, we discuss the impacts, limitations, sim-to-real generalization, and future work of MetaUrban.